Jordan is tracking a recent online purchase. The shipping costs state that the item will be shipped in a 24-inch long box with a volume of 2,880 cubic inches. The width of the box is seven inches less than the height.
Answer by ankor@dixie-net.com(22740)

Jordan is tracking a recent online purchase. The shipping costs state that the item will be shipped in a 24-inch long box with a volume of 2,880 cubic inches. The width of the box is seven inches less than the height.
:
Assume you want the dimensions of the box.
:
let h = the height of the box
then
(h-7) = the width of the box
Given the length = 24
The volume equation L*w*h
24*(h-7)*h = 2880
24(h^2 - 7h) = 2880
divide both sides by 24
h^2 - 7h = 120
A quadratic equation
h^2 - 7h - 120 = 0
Factors to
(h-15)(h+8) = 0
the positive solution is all we want here
h = 15 inches is the height
and obviously,
17-7 = 8 in is the width
:
:
:
Check: 24 * 8 * 15 = 2880